Be sure to answer all parts. Calculate the pH of a solution that is: 1.6 x 10-2 M in HCI. Reference
Answer #1

PH= -log[H+]

HCl(aq) -------> H+(aq) + Cl-(aq)

Given: [H+] = 1.6 x 10-2M

PH= -log[H+] = -log[1.6 x 10-2] = -log(1.6) - log(10-2) [since , log(axb) = loga + logb and logab = b loga]

=-0.204-(-2log10) [ log10=1]

  PH =2-0.204= 1.795